# Spin-orbit interaction in tubular prismatic nanowires

### v1

@@ III.3.3 Tri-Hex shell
Figure 11: Energy dispersions for the corner states of the 8-nm-wide Tri-Hex shells grown on different cores: $R_{\mathrm{c}}=40$ nm (a), $R_{\mathrm{c}}=20$ nm (b). The inset in panel (b) shows a zoomed-in view of the states around $k_{z}=0$. The energies are shifted with respect to the ground energy at $k_{z}=0$.
@@
The transverse energies [Fig. 3(c)] and the low-energy electron localization [Fig. 3(c)] of the Tri-Hex shells resemble those of the triangular ones. Since the $\Delta_{\mathrm{Tri-Hex}}$ gap can be comparable to or even larger than the corresponding value for triangular wires $\Delta_{\mathrm{Tri}}$ [57], we focus on the six lowest states localized in the corner regions.
@@
The corner areas of Tri-Hex structures are much larger than those of hexagonal and triangular shells grown on cores with the same radius ($R_{\mathrm{c}}$) (Fig. 1). When $r=2$ nm, the area where the SOI coefficients (Eq. 21) are nonzero is the same as in the case of hexagonal wires. However, this area constitutes a very small fraction of the Tri-Hex cross section, so the impact of SOI is minimal [Fig. 11(a)]. Further, we increased the confinement by reducing the core radius by half while keeping $r=2$ nm and $d=8$ nm. This change increases the energy separation of the lowest six corner states at $k_{z}=0$ from 31 to 76 meV, and amplifies the effects due to SOI [Fig. 11(b)]. Since at the same time the ratio $d/R_{\mathrm{s}}$ increases, the energy splitting between the ground state and the first excited state at $k_{z}=0$ also increases [inset to Fig. 11(b)] and the corner localization softens, but the probability maxima remain well separated.
@@
Figure 12: Energy dispersions for the corner states of different Tri-Hex shells, with the intermixing length equal to 75% of the corner area’s height. The energies are shifted with respect to the ground energy at $k_{z}=0$.
[…]
Interestingly, when the core material molecules penetrate deep into the shell, the effects due to SOI are relatively robust to the increase in shell width. If $R_{\mathrm{c}}=40$ nm, then increasing $d$ from 8 to 24 nm reduces the impact of SOI, but it remains significant [Fig. 12(b)]. For the smaller shell, increasing $d$ by 50% only weakly influences the energy dispersions [Figs. 12(d)]. Due to its geometry, an increase in shell width results in a smaller relative enlargement of the Tri-Hex corner areas compared to triangular and hexagonal shell structures. As a result, the change in the effect of SOI is also smaller.
@@
Figure 13: Energy dispersions for the corner states of the 8-nm-wide Tri-Hex shell in the presence of an external electric field. Here, $R_{\mathrm{c}}=20$ nm, $r=d=8$ nm, $E=0.1$ mV/nm, $\varphi=\pi/6$. The energies are shifted with respect to the ground energy at $k_{z}=0$.
@@ IV Conclusions
In summary, we have studied the impact of the SOI due to the extended band offset potential and an external electric field on the energies of hexagonal, triangular, and Tri-Hex tubular wires, which are the outer layers of core-shell nanowires. Our results are derived with the $\bm{k}\!\cdot\!\bm{p}$ method for the shell of a core-shell nanowire, and are compatible to calculations performed by other authors [49, 39] for the core. The presence of SOI in the shell has been experimentally observed, for example in o recent study of magnetoconductance oscillations [60], and to the best of our knowledge, band-structure based models of SOI in prismatic shells are missing or are rare in the literature.
We have shown that, if the SOI is caused only by the conduction-band offset along the heterojunction, then the effect of SOI on the energies of hexagonal shells is qualitatively different from the impact on the spectra of triangular and Tri-Hex structures. For all three geometries studied (hexagonal, triangular, and combined), the SOI due to the band offset potential lifts only the 4-fold degeneracies at finite $k_{z}$, but in the case of hexagonal tubes, it splits the states into two 2-fold degenerate states such that the states of hexagonal shells become pairwise degenerate (though spin is no longer conserved). In contrast, the 4-fold degenerate levels of triangular and Tri-Hex tubes split into three states: the lowest and highest states are nondegenerate, while the middle state is 2-fold degenerate.
Qualitatively, the SOI due to the extended band offset potential affects the corner- and side-localized states in the same way, but the impact on the lower-energy states is stronger. If the intermixing length is kept short and constant while the shell thickness increases, i.e. as clean shells are grown, the impact of SOI on the energy states decreases together with the cross-sectional fraction, where the SOI coefficients are nonzero. However, when the shell thickness increases the SOI has a much stronger effect on the energies of triangular tubes compared to the hexagonal ones.
The SOI in polygonal tubes can be enhanced or controlled through engineering of the band offset potential. Specifically, if a tilted potential that penetrates deep into the shell is created, i.e. extended band offset potential is present, the SOI coefficients become nonzero over a significant fraction of the cross-sectional area. This leads to a considerable enhancement of SOI which, according to our results, for triangular and hexagonal wires is maximized for $r=d$, although for Tri-Hex structures the strongest effects occur for $r<d$. Such SOI engineering is advantageous, for example, in the context of Majorana states. Additionally, the extended band offset potential increases the confinement, thereby amplifying characteristics associated with polygonal cross sections, such as electron accumulation along the sharp edges and the energetic separation of the lowest corner states.

### v2

@@ III.3.3 Tri-Hex shell
Figure 11: Energy dispersions for the corner states of the 8-nm-wide Tri-Hex shells grown on different cores: (a) $R_{\mathrm{c}}=40$ nm, (b) $R_{\mathrm{c}}=20$ nm. The inset in (b) shows a close-up view of the states around $k_{z}=0$. The energies are shifted with respect to the ground state energy at $k_{z}=0$.
@@
The transverse energies [Fig. 3(c)] and the low-energy electron localization [Fig. 3(c)] of the Tri-Hex shells resemble those of the triangular ones. Since the $\Delta_{\mathrm{Tri\textrm{-}Hex}}$ gap can be comparable to or even larger than the corresponding value for triangular wires ($\Delta_{\mathrm{Tri}}$) [57], and the states above $\Delta_{\mathrm{Tri\textrm{-}Hex}}$ are localized in the corner regions, we focus on the six lowest states localized in the corners.
@@
The corner areas of Tri-Hex structures are much larger than those of hexagonal and triangular shells grown on cores with the same radius (Fig. 1). When $r=2$ nm, the area where the SOI coefficients [Eq. (21)] are nonzero is the same as in the case of hexagonal wires. However, this area constitutes a very small fraction of the Tri-Hex cross section, so the impact of SOI is minimal [Fig. 11(a)]. Further, we increased this fraction by reducing the core radius by half while keeping $r=2$ nm and $d=8$ nm, which resulted in the amplification of the SOI [Fig. 11(b)]. Moreover, the energy splitting between the ground state and the first excited state at $k_{z}=0$ increased [inset to Fig. 11(b)], along with the ratio $d/R_{\mathrm{s}}$.
@@
Figure 12: Energy dispersions for the corner states of different Tri-Hex shells, with the intermixing length equal to 75% of the corner area’s height. The energies are shifted with respect to the ground state energy at $k_{z}=0$.
@@
Figure 13: Energy dispersions for the corner states of the 8-nm-wide Tri-Hex shell in the presence of an external electric field. Here, $R_{\mathrm{c}}=20$ nm, $r=d=8$ nm, $E=0.1$ mV/nm, $\varphi=\pi/6$. The energies are shifted with respect to the ground state energy at $k_{z}=0$.
As in the case of hexagonal and triangular shells, we increase the parameter $r$, representing the intermixing length of the core and shell materials, to 75% of the distance between the core-shell interface and the shell external vertices. The enhancement due to the extended band-offset potential is considerably stronger than for the triangular wires. The energy dispersions for the 8-nm-wide shell [Figs. 12(a) and 12(c)], grown either on a core inscribed in a 40- or 20-nm circle, indicate that both structures behave as three independent and identical wires. The absence of interaction is also confirmed by the crossings of energy states at finite $k_{z}$ in the presence of a perpendicular electric field, when the three effective wires are no longer identical (Fig. 13). Moreover, the strength of SOI is of the order of that obtained for the triangular shells, i.e., $\alpha$ varies between 4 and 28 $\mathrm{meV}\hskip 0.85355pt\mathrm{nm}$.
Interestingly, when the core material penetrates deep into the shell, the effects due to SOI are relatively robust to the increase in shell width. If $R_{\mathrm{c}}=40$ nm, then increasing $d$ from 8 to 24 nm reduces the impact of SOI, but it remains significant [Fig. 12(b)]. For the smaller shell, increasing $d$ by 50% only weakly influences the energy dispersions [Fig. 12(d)]. Due to its geometry, an increase in shell width results in a smaller relative enlargement of the Tri-Hex corner areas compared to triangular and hexagonal shell structures. As a result, the change in the effect of SOI is also smaller.
@@ IV Conclusions
In summary, we studied the impact of the SOI due to the extended band-offset potential and an external electric field on the energies of hexagonal, triangular, and Tri-Hex tubular wires, which are the outer layers of core-shell nanowires. Our results are derived with the $\bm{k}\!\cdot\!\bm{p}$ method for the shell and are compatible to calculations performed by other authors [49, 39, 60] for the core. The presence of SOI in the shell has been experimentally observed, for example, in a recent study of magnetoconductance oscillations [61], and to the best of our knowledge, band-structure-based models of SOI in prismatic shells are missing or are rare in the literature.
We show that, if the SOI is caused only by the conduction-band offset along the heterojunction, then the effect of SOI on the energies of hexagonal shells is qualitatively different from the impact on the spectra of triangular and Tri-Hex structures. For all three geometries studied (hexagonal, triangular, and combined), the SOI due to the band-offset potential lifts only the fourfold degeneracies at finite $k_{z}$, but in the case of hexagonal tubes, it splits the states into two twofold degenerate states such that the states of hexagonal shells become pairwise degenerate (though spin is no longer conserved). In contrast, the fourfold degenerate levels of triangular and Tri-Hex tubes split into three states: The lowest and highest states are nondegenerate, while the middle state is twofold degenerate.
Qualitatively, the SOI due to the extended band-offset potential affects the corner- and side-localized states in the same way, but the impact on the lower-energy states is stronger. If the intermixing length is kept short and constant while the shell thickness increases, i.e., as clean shells are grown, the impact of SOI on the energy states decreases together with the cross-sectional fraction, where the SOI coefficients are nonzero. However, when the shell thickness increases the SOI has a much stronger effect on the energies of the triangular tubes compared to the hexagonal ones.
The SOI in polygonal tubes can be enhanced or controlled through engineering of the band-offset potential. Specifically, if a tilted potential that penetrates deep into the shell is created, i.e., an extended band-offset potential is present, the SOI coefficients become nonzero over a significant fraction of the cross-sectional area. This leads to a considerable enhancement of SOI which, according to our results, for triangular and hexagonal wires is maximized for $r=d$, although for Tri-Hex structures the strongest effects occur for $r<d$. Such SOI engineering is advantageous, for example, in the context of Majorana states. Additionally, the extended band-offset potential increases the confinement, thereby amplifying characteristics associated with polygonal cross sections, such as electron accumulation along the sharp edges and the energetic separation of the lowest corner states.
@@
The data supporting the figures of this paper are available at the Zenodo repository [62].